# MT-HCCAR: Multi-Task Deep Learning with Hierarchical Classification and Attention-based Regression for Cloud Property Retrieval

### v1

@@ V-A Experiment setting
Besides the above two baseline methods, four ablation study models are also used to evaluate the effectiveness of different modules in our model. In these four models, specific modules are systematically removed from our proposed model, MT-HCCAR, allowing us to assess the impact of these modifications on model performance.
  • SEQ: A sequential structure model consists of three separate networks. The first one classifies [cloudy, cloud-free], the second one classifies [liquid-cloud, ice-cloud], and the third network makes regressive predictions on COT for cloudy pixels, respectively.
  • MT-CR: MTL method combines the classification network and regression network in an end-to-end model.
  • MT-HCR: Multi-task learning model with the HC network classification task and a regression network in the multi-task learning approach.
  • MT-HCCR: Multi-task learning model with hierarchical classification and classification-assisted regression. Only the cross-attention module is removed from MT-HCCAR.
After comparing different models on the same dataset OCI, we further use 10-fold cross-validation [43, 44] on four MTL-based models MT-CR, MT-HCR, MT-HCCR, and MT-HCCAR to assess the usefulness of modules and compare the difference in model performance between our models and two existing methods for cloud property retrieval. Additionally, each of the four models will be trained by data from three sensors (OCI, VIIRS, and ABI) respectively.
V-B Evaluation metrics
@@
We use three types of metrics to evaluate our work: 1) metrics for the classification task, 2) metrics for the regression task, 3) metrics from an Earth science perspective.
V-B1 Metrics for classification performance
We use Accuracy and Average precision to evaluate classification performance.
[…]
$Accuracy_{i}=(TP_{i}+TN_{i})/(TP_{i}+TN_{i}+FP_{i}+FN_{i})$ (12)
, where $TP_{i}$, $TN_{i}$, $FP_{i}$, and $FN_{i}$ are the number of true positives, true negatives, false positives, and false negatives for class $i$, respectively.
Area under precision-recall curve ($AU(\overline{PRC})_{w}$, $AU(\overline{PRC})_{c}$): Weighted area under the precision-recall curve ($AU(\overline{PRC}_{w}$)) is the weighted mean of precisions at each threshold $h$ for class $i$, with the weight as substract between recall at threshold $h$ and the recall at threshold $h-1$. The area under the precision-recall curve for each label $AU(\overline{PRC})_{c}$ is calculated for each binary label in $l=\{C,\bar{C},CL,CI\}$. For the weighted area under the precision-recall curve, the recall and the precision at threshold $h$ are:
$R^{h}_{w}=\frac{\sum_{l}TP^{h}_{c}}{\sum_{l}TP^{h}_{c}+\sum_{l}FN^{h}_{c}},P^{h}_{w}=\frac{\sum_{l}TP^{h}_{c}}{\sum_{l}TP^{h}_{c}+\sum_{l}FP^{h}_{c}}$ (13)
And for the area under the precision-recall curve of each label $c$, the recall and the precision at threshold $h$ are:
$R^{h}_{c}=\frac{TP^{h}_{c}}{TP^{h}_{c}+FN^{h}_{c}},P^{h}_{c}=\frac{TP^{h}_{c}}{TP^{h}_{c}+FP^{h}_{c}}$ (14)
With Equation 13 and Equation 14, $AU(\overline{PRC})_{w}$ and $AU(\overline{PRC})_{c}$ are:
$AU(\overline{PRC})_{w}=\sum_{h}(R^{h}_{w}-R^{h-1}_{w})P^{h}_{w}$ (15)
$AU(\overline{PRC})_{c}=\sum_{h}(R^{h}_{c}-R^{h-1}_{c})P^{h}_{c}$ (16)
V-B2 Metrics for regression performance
We use MSE and $R^{2}$ to evaluate regression performance.
Mean squared error (MSE) is the average squared difference between the true and predicted COT values. We calculate MSE across all classes and MSE for each class of [cloud-free, liquid-phase, ice-phase].
$MSE=\sum_{i=0}^{n}(y_{i}^{cot}-\hat{y}_{i}^{cot})$ (17)
Coefficient of determination ($R^{2}$) measures how close the predicted value is to the true value. $R^{2}$ across all classes and $R^{2}$ for each class.
$R^{2}=1-\frac{\sum_{i=0}^{n}(y_{i}^{cot}-\hat{y}_{i}^{cot})^{2}}{\sum_{i=0}^{n}(y_{i}^{cot}-\overline{y}^{cot})^{2}}$ (18)
, where $\overline{y}^{cot}$ is the mean value of all true COT values $[y_{1}^{reg},y_{2}^{reg},\ldots,y_{n}^{reg}]$.
V-B3 Earth science metrics
Besides the above commonly used data mining metrics, we also evaluate how a machine learning model could be used for actual satellite missions.
@@ V-B3 Earth science metrics
Fraction of pixels meeting PACE goals (FMG) is an evaluation metric defined by the PACE Validation Plan [45] based on scientific requirements and expectations. For each pixel, the relative error $Error_{i}=|\frac{y^{(cot)_{i}}-\hat{y}^{(cot)_{i}}}{y^{(cot)_{i}}}|$. Then fraction meeting goals represents the percentage of pixels whose relative error is less than 0.25 (for liquid clouds) or 0.35 (for ice clouds).
The Validation Plan defines a predictive model with satisfactory performance as one where this goal is met for 65% of cloudy pixels. Note that this mission goal only applies to cases where the true original COT value is 5 or more (log10 COT > 0.7).
@@
TABLE I: Model performance Comparison with Baseline Models and Ablation Study - classification results. Models are trained by OCI.
[…]
@@
TABLE II: Model performance Comparison with Baseline Models and Ablation Study - regression results. Models are trained by OCI.
@@
V-C Comparison with baseline models
[…]
Turning attention to the Cloud Optical Thickness (COT) regression task, the Multi-Layer Perceptron (MLP)-based baseline method introduced by [42] yields results comparable to the SEQ model, but much worse than MTL-based models. This observation underscores the efficiency of MTL-based models for of COT regression.
V-D Ablation study
Figure 3: Scatter plot with PDF distribution of true COT values (y-axis) and predicted COT values (x-axis).
The outcomes obtained from models SEQ through MT-HCCAR in Table I and Table II underscore the efficacy of our MTL structure and the incorporated modules, namely Hierarchical Classification (HC) and Cross-Attention Residual (CAR), in enhancing both classification and regression performances. A comparative analysis reveals notable improvements across all metrics when contrasting the results of models through SEQ to MT-HCCAR. Specifically, when compared to the MT-CR model, the introduction of hierarchical classification MT-HCR demonstrates superior accuracy in binary classification ([cloudy, cloud-free]) and further refines the classification of liquid and ice cloud phases, thereby contributing to enhancements in all regression metrics. The performance of the model does not see any drop from MT-HCR to MT-HCCAR, even though the models focus on improving regression by using CAR.
Moreover, the efficacy of the CAR module is validated in the comparison between MT-HCCAR and MT-HCR, showcasing its utility in improving both classification and regression outcomes. Furthermore, the adoption of the cross-attention module is substantiated by the superior performance observed in MT-HCCR compared to MT-HCCAR, underscoring the module’s effectiveness in facilitating information exchange between hierarchical classification and regression tasks.
@@ V-D Ablation study
In order to visually depict the enhancements facilitated by the HC module and the CAR module, Figure 3 presents a scatter plot illustrating the distribution of ground truth Cloud Optical Thickness (COT) values and the corresponding predicted COT values for all instances within the test set. The integration of the HC module and CAR module in our model yields more accurate predictions compared to MT-CR and MT-HCR, as evidenced by two notable observations in the figure. Firstly, the Probability Density Function (PDF) distribution of predictions closely aligns with the true distribution, indicating an improved fidelity to the actual COT values. Secondly, the scatter points exhibit a greater concentration along the diagonal line, further corroborating the model’s enhanced precision when incorporating the HC and CAR modules.
V-E Earth science evaluation
V-E1 Applicability for satellite mission
@@
Satellite-derived cloud masks are generally evaluated either through comparison against a human-labeled image set, or against ground-based, airborne, or spaceborne observations from sensors that have much greater sensitivity for cloud detection than the image in question (typically active sensors such as radar and/or lidar). Accuracies tend to depend on surface type and illumination conditions; for daytime scenes such as those simulated here, some reported accuracies for NASA’s widely-used MODIS cloud mask are 0.850 (Ackerman et al.[46]), 0.778 (Wang et al. [47]), and 0.894 (Kotarba [48]). Accuracy tends to be worse than average in polar regions and better for non-polar land and ocean [49]. Most commonly it is optically thin clouds (e.g. subvisual cirrus) that are missed. For instance, Ackerman et al. [46] found that 90% of misclassified pixels in MODIS over a vegetated ground site in the USA had a COT under 0.4. Chen et al. [42] developed an NN cloud mask for MODIS, also based on RT simulations, and found an accuracy of 0.985 on training data but lower accuracies in the range of 0.739-0.885 on real data depending on region and time of year. Although real-world conditions are more complex than a simulation, and each sensor’s measurement characteristics result in different sensitivity to clouds, the results suggest our models’ accuracies are competitive with current approaches. Note, that the cloud phase is less commonly validated in this way.
@@
Obtaining a validation-grade COT measurement is exceptionally challenging, and reference datasets for COT lack the comprehensive coverage available for other cloud properties like cloud mask and cloud height[50]. Consequently, existing comparisons primarily focus on assessing agreement levels among different remotely sensed datasets. As previously mentioned, the PACE mission aims to achieve an error rate of less than 25% (for liquid clouds) or 35% (for ice clouds) in determining COT for clouds with true COT > 5, at least 65% of the time. Our network’s performance indicates its potential to meet these objectives. Notably, the larger allowable uncertainty for ice clouds in the PACE mission is rooted in the expectation that accurate determination would be more challenging due to significant variations in ice crystal shape, size, and roughness in real-world scenarios. In contrast, our results demonstrate that the predicted uncertainty for ice COT can be smaller than that for liquid COT. Conventional approaches often assume specific ice crystal shapes and roughness, leading to increased uncertainty in retrieved ice COT. Our findings suggest that satellite measurements inherently contain information related to these parameters, which the networks are adept at learning.
V-E2 Performance of different satellite sensors
@@
The classification tasks exhibit high accuracies across all sensors, with virtually indistinguishable performance from a scientific perspective. This uniformity is unsurprising, despite the considerable variance in the number of bands among sensors. The fundamental nature of cloud masking in Earth science prompts many sensors to incorporate a common set of bands proven to be effective for this purpose, with additional bands often designed for diverse applications. For instance, OCI’s hyperspectral bands support the measurement of different ocean plankton species [3], revealing subtle spectral differences not readily discernible in multispectral data. In the regression task, OCI and VIIRS outperform ABI, indicating the utility of additional bands in predicting COT. The unexpectedly slightly superior performance of VIIRS over OCI may suggest that OCI’s substantially larger feature space makes finding an optimal solution during training more challenging or could be attributed to stochastic variation. Adding training epochs for OCI may lead to better results.
VI Strategies for Selecting the Optimal Model

### v2

@@ 5 Experiments
Besides the above two baseline methods, four ablation study models are also used to evaluate the effectiveness of different modules in our model. The comparison of the four models and our MT-HCCAR is shown in Table 1. The detailed differences can be found in the supplementary material.
Table 1: Comparison of models in the ablation study. STL: single-task learning.
Model Structure Classification Regression
SEQ STL Regular Regular
MT-CR MTL Regular Regular
MT-HCR MTL HC Regular
MT-HCCR MTL HC CAR without Attn.
MT-HCCAR MTL HC CAR
5.1 Experiment Setup
Our model was implemented using the Python deep-learning library PyTorch. All baseline models and proposed models were tested on a single GPU on Kaggle. The dataset comprises satellite data simulation of instruments OCI, VIIRS, and ABI. The three datasets encompassing $N=250,000$ instances.
We split data into 62.5%, 22.5%, and 10.0% for training, validation, and test sets, respectively. All models are trained with the same hyperparameters including learning rate = $1e^{-5}$, batch size = $64$, and training epochs = $500$. All experiment results are mean values of metrics from a 10-fold cross-validation.
Simulations of OCI, VIIRS, and ABI are utilized instead of observations as our labeled training dataset principally because there is no suitable comprehensive reference truth dataset. While the standard retrieval products from these sensors could be used to train networks, for example, these products have known limitations so there would be a risk of the model training to the artifacts in these products. In a similar vein, comparing the results of applying the trained NN to real satellite observations of these standard products will be instructive to get a general sense of reasonableness but not to gauge their absolute performance. It is expected that, by using a more realistic training set of simulations than was used to develop the physically-based retrievals (which include many simplifying assumptions out of the computational necessities from decades ago when these approaches were developed), the NN model should outperform them.
5.2 Evaluation Metrics
@@
We use three types of metrics to evaluate our work, including metrics for the classification task, metrics for the regression task, and metrics from an Earth science perspective.
Metrics for classification performance. We use Accuracy and Average precision to evaluate classification performance. Cloud masking accuracy (ACCbi) is the fraction of correct predictions of two big categories [cloudy, cloud-free] by our model. The area under the precision-recall curve ($AU(\overline{PRC})_{c}$) is calculated for each label. The weighted area under the precision-recall curve ($AU(\overline{PRC})_{w}$) is the weighted mean of precisions across all labels at each threshold $h$ for class $i$, with the weight as subtracted between recall at threshold $h$ and the recall at threshold $h-1$, where the number of thresholds is close to infinity. The weighted precision and recall are: $R^{h}_{w}=\frac{\sum_{l}TP^{h}_{c}}{\sum_{l}TP^{h}_{c}+\sum_{l}FN^{h}_{c}}$ and $P^{h}_{w}=\frac{\sum_{l}TP^{h}_{c}}{\sum_{l}TP^{h}_{c}+\sum_{l}FP^{h}_{c}}$, then $AU(\overline{PRC})_{w}$ is: $AU(\overline{PRC})_{w}=\sum_{h}(R^{h}_{w}-R^{h-1}_{w})P^{h}_{w}$.
Metrics for regression performance. We use MSE and $R^{2}$ to evaluate regression performance. Mean squared error (MSE) is the average squared difference between the true and predicted COT values. $MSE=\sum_{i=0}^{n}(y_{i}^{cot}-\hat{y}_{i}^{cot})$. The coefficient of determination ($R^{2}$) measures how close the predicted value is to the true value. $R^{2}=1-\frac{\sum_{i=0}^{n}(y_{i}^{cot}-\hat{y}_{i}^{cot})^{2}}{\sum_{i=0}^{n}(y_{i}^{cot}-\overline{y}^{cot})^{2}}$, where $\overline{y}^{cot}$ is the mean value of all true COT values $[y_{1}^{reg},y_{2}^{reg},\ldots,y_{n}^{reg}]$.
@@
Earth science metrics. Besides the above metrics, we also evaluate how a machine learning model could be used for actual satellite missions. The fraction of pixels meeting PACE goals (FMG) is an evaluation metric defined by the PACE Validation Plan [4] based on scientific requirements and expectations. For each pixel, the relative error $Error_{i}=|\frac{y^{(cot)_{i}}-\hat{y}^{(cot)_{i}}}{y^{(cot)_{i}}}|$. Then FMG represents the percentage of pixels whose relative error is less than 0.25 (for liquid clouds) or 0.35 (for ice clouds). The Validation Plan defines a predictive model with satisfactory performance as one where this goal is met for 65% of cloudy pixels. Note that this mission goal only applies to cases where the true original COT value is 5 or more (log10 COT > 0.7).
5.3 Comparison with Baseline Models
Tables 2 and 3 present a comprehensive evaluation of our proposed model MT-HCCAR against two baseline methods, demonstrating superior performance across both classification and regression metrics. Specifically, in the assessment of classification tasks, ACCbi pertains to cloud masking performance, while $AU(\overline{PRC})_{w}$ and $AU(\overline{PRC})_{c}$ quantify the performance of both cloud masking and cloud phase classification.
Upon close examination of Table 2, the RF classifier by [17] attains superior ACCbi, $AU(\overline{PRC})_{w}$, and $AU(\overline{PRC})_{c}$ in comparison to the SEQ model and the simplest MTL based model, MT-CR, within the ablation study. However, MTL-based models featuring the HC module including MT-HCR, MT-HCCR, and MT-HCCAR, surpass the performance of the RF classifier in these metrics. Shifting the focus to the results of the COT regression task in Table 3, the MLP-based baseline method introduced by [8] produces results on par with the SEQ model but significantly inferior to MTL-based models. This finding emphasizes the effectiveness of MTL-based models for COT regression.
@@ 5.3 Comparison with Baseline Models
Table 2: Model performance comparison with baseline models and ablation study - classification results. Models are trained on OCI.
Model ACCbi $AU(\overline{PRC})_{w}$ $AU(\overline{PRC})_{c}$
@@
Table 3: Model performance comparison with baseline models and ablation study - regression results. Models are trained on OCI. The model [17] is not applicable for the regression task.
@@
Table 4: Model performance for VIIRS and ABI datasets.
Dataset Model ACCbi $AU(PRC)_{w}$ MSE $R^{2}$ FMG ($CL$, $CI$)
VIIRS MT-CR 0.976 0.939 0.043 0.810 48.52% 72.73%
MT-HCR 0.989 0.996 0.030 0.874 66.23% 82.62%
MT-HCCAR 0.989 0.997 0.025 0.897 68.20% 82.69%
ABI MT-CR 0.973 0.969 0.057 0.751 47.87% 71.81%
MT-HCR 0.986 0.996 0.038 0.840 66.58% 82.55%
MT-HCCAR 0.987 0.996 0.032 0.866 68.54% 83.34%
5.4 Ablation Study
The ablation study with five models from SEQ to MT-HCCAR for OCI, as depicted in the lower part of Tables 2 and 3, highlights the effectiveness of our model in improving both classification and regression performance. The comparative analysis demonstrates significant improvements across all metrics from SEQ to MT-CR, which shows the usefulness of the MTL structure. For MTL methods, MT-HCR outperforms MT-CR in binary classification (cloudy, cloud-free) and refines liquid and ice cloud phase classification, leading to enhancements in regression metrics. This proves the effectiveness of the HC module. Moreover, despite focusing on regression, the introduction of CAR in MT-HCCAR maintains the performance of classification tasks. The efficacy of the CAR module is further validated by comparing MT-HCCAR to MT-HCR, indicating improvements in all metrics. Additionally, the adoption of the cross-attention module is substantiated by superior performance in MT-HCCR compared to MT-HCCAR, emphasizing its role in facilitating information exchange between hierarchical classification and regression.
To confirm the generalizability of our proposed model, we did a further ablation study with VIIRS and ABI. As shown in Table 4, the application of MTL-based models reveals a similar trend of performance enhancement when integrating HC and CAR modules. Notably, MT-HCCAR achieves the highest performance across both classification and regression tasks among the MTL models across OCI, VIIRS, and ABI datasets, underscoring the generalization capabilities of the introduced model MT-HCCAR. The detailed results of cross-validation and model selection are in the supplementary material.
@@ 5.4 Ablation Study
To visually illustrate the improvements facilitated by the HC and CAR modules, Figure 3 presents a scatter plot showing the distribution of true COT values against predicted COT values for all instances in the test set. Integration of HC and CAR modules results in more accurate predictions compared to MT-CR and MT-HCR. Notably, the predicted Probability Density Function (PDF) aligns closely with the true distribution, indicating improved fidelity to actual COT values. Additionally, scatter points cluster more tightly along the diagonal line, confirming the model’s enhanced precision when HC and CAR modules are incorporated into the model.
Figure 3: Scatter plot with PDF distribution of true COT values (y-axis) and predicted COT values (x-axis) for the OCI dataset.
5.5 Earth Science Evaluation
@@
Quality of performance. Satellite cloud masks are generally evaluated either through comparison against ground-based, airborne, or spaceborne observations from radar/lidar sensors that have much greater sensitivity for cloud detection than imagers in question. Accuracies tend to depend on the surface type and illumination conditions; for daytime scenes such as those simulated here, reported accuracies for NASA’s widely-used MODIS cloud mask are 0.850, 0.778, and 0.894. Chen et al. [8] developed an NN cloud mask for MODIS based on RT simulations and found an accuracy of 0.985 on training data but lower accuracies in the range of 0.739-0.885 on real data depending on region and time of year. Although real-world conditions are more complex than a simulation, the results suggest our models are competitive with current approaches. Note, that the cloud phase is less commonly validated in this way.
@@
Obtaining a validation-grade measurement of COT presents significant challenges, with reference datasets lacking comprehensive coverage [19]. Consequently, existing comparisons primarily focus on evaluating agreement among different remotely sensed datasets. Our network’s performance suggests its potential to fulfill the objectives outlined in the Validation Plan [4]. Notably, PACE’s allowance for larger uncertainty in ice cloud measurements stems from the anticipated difficulties arising from variations in ice crystal properties in real-world scenarios. However, our results show that the uncertainty in ice COT can be smaller than that for liquid COT. Traditional methods often assume specific ice crystal properties, resulting in increased uncertainty in retrieved ice COT. Our findings indicate that satellite measurements inherently contain information about these properties, which neural networks are proficient at learning.
@@
Performance of different satellite sensors. We also compared how the models perform for different satellite sensors based on the results in Tables 23 and 4. The classification tasks exhibit high accuracy across all sensors, with virtually indistinguishable performance from a scientific perspective. The fundamental nature of cloud masking in Earth science prompts many sensors to incorporate a common set of bands proven to be effective for this purpose, with additional bands often designed for diverse applications. For instance, OCI’s hyperspectral bands support the measurement of different ocean plankton species [28], revealing subtle spectral differences not readily discernible in multispectral data. In the regression task, OCI and VIIRS outperform ABI, indicating the utility of additional bands in predicting COT. The unexpectedly slightly superior performance of VIIRS over OCI may suggest that OCI’s substantially larger feature space makes finding an optimal solution during training more challenging or could be attributed to stochastic variation. Adding training epochs for OCI may lead to better results, considering the number of bands.
6 Conclusions
In this study, we present MT-HCCAR, an end-to-end multi-task learning model tailored for cloud property retrieval on a simulated OCI satellite dataset in the PACE project, tackling tasks including cloud masking, cloud phase classification, and COT prediction. The model is implemented on three sensors’ simulated datasets (OCI, VIIRS, and ABI), respectively, to examine its generalization. Comparative analyses against two baseline methods and ablation studies underscore the effectiveness of the HC module and the CAR module, enhancing performance in both classification and regression tasks. The ablation study establishes MT-HCCAR’s superior performance across different datasets and multiple evaluation metrics. The positive results affirm our model’s capability to address real-world challenges in cloud property retrieval and other multi-task applications. Future research endeavors will involve applying the model to spatial or temporal OCI images, co-located in space and time with VIIRS and ABI post-PACE launch, to assess its performance, and consistency, and enable detailed comparisons with deep learning models and non-machine-learning approaches.
{credits}
6.0.1 Acknowledgements
Appendix 0.A Experiment Settings of Ablation Study
Four ablation study models are also used to evaluate the effectiveness of different modules in our model. The four models are 1) SEQ: A sequential structure model consists of three separate networks. The first one classifies [cloudy, cloud-free], the second one classifies [liquid-cloud, ice-cloud], and the third network makes regressive predictions on COT for cloudy pixels, respectively. 2) MT-CR: The MTL method combines the classification network and regression network in an end-to-end model. 3) MT-HCR: MTL model with the HC network classification task and a regression network in the MTL approach. 4) MT-HCCR: MTL model with hierarchical classification and classification-assisted regression, where only the cross-attention module is omitted from MT-HCCAR.
The purpose of comparing the four models with MT-HCCAR is to prove the effectiveness of multi-task learning (MTL), the effectiveness of the hierarchical classification sub-network (HC), the effectiveness of the auxiliary classification module in the CAR sub-network, and the effectiveness of the cross-attention module in the CAR sub-network, respectively. As described in the main manuscript, the objectives of this work include 1) classification of cloud masks (cloud existence), 2) classification of ice-cloud and liquid-cloud for cloudy pixels, and 3) regression of COT for cloudy pixels. Therefore, both classification networks and regression networks are necessary for whatever model we use to complete the three tasks. The difference lies in how we combine the networks, and what exact method we use to improve the performance of each task.
To provide a better understanding of the differences in models compared to the ablation study, below are figures of model architectures of each model. Figure 4 to Figure 7 illustrate model structures of SEQ, MT-CR, MT-HCCR, and MT-HCCR, respectively.
Figure 4: The model SEQ, where SEQ is the abbreviation of Sequential Model. MTL method is not used for SEQ.
Figure 5: The model MT-CR, where MT is the abbreviation of MTL, C represents Classification and R represents Regression. MTL method is used, but neither HC nor CAR sub-networks are used.
Figure 6: The model MT-HCR, where HC represents the proposed HC sub-network and R represents Regression. MTL method and HC are used, while the CAR sub-network is not.
Figure 7: The model MT-HCCR, where HC represents the proposed HC sub-network, CR represents the CAR model without the cross-attention module (A).
Appendix 0.B Cross-Validation and Model Selection